BI/TB on AD CCC (follow-up), future use of CALIFES and VFC status

Europe/Zurich
866/2-D05 (CERN)

866/2-D05

CERN

30
Show room on map

Actions:

AD CCC:

  1. Offset jumps remain at injection (Miguel)
  2. Issues with calibration generators (Miguel)
  3. Operational integration (Diogo)

LHC Status:

  1. BSRT software and electronics (digital cameras) upgrade to be decided (Enrico/Diogo/L. Jensen)
  2. Diamond BLM integration with OP-LHC (Stephane/Enrico)
  3. BPM capture data sometimes missing (diagnostics to be improved) (L. Jensen/Guillaume/Manfred/BE-CO)
  4. "Dancing BPMs" (Manfred)
  5. MD blocks #4 and #5 to be planned LBOC meeting (Enrico)

VFC update:

  1. Preparing for tendering process (expected mid-Sept 2016) (Thibaut/Andrea)
  2. 50 VFCs ordered from Norcott - see last BI-TB for who gets them (expected early Oct)
  3. DDR3 access issue remaining - need help from Altera (Hermann/Thibaut/Rhodri/Andrea/Christos)
There are minutes attached to this event. Show them.